OpenStudy (anonymous):

The diagonal of a computer monitor is sqrt(250) inches long. Which expression has the same value as sqrt(250)? A. 5sqrt(5) B. 5sqrt(10) C. 10sqrt(5) D. 25sqrt(10) I think that its D????

OpenStudy (studygurl14):

D is incorrect. It is A

OpenStudy (studygurl14):

Because... \(\large\sqrt{250}\rightarrow\sqrt{25\times10}\rightarrow\sqrt{25}\times\sqrt{10}\rightarrow5\sqrt{10}\)

OpenStudy (anonymous):

ohhhhhh ok THANKYOU!! :D

OpenStudy (studygurl14):

\(\tt\large\color{#008AAB}{You're~Welcome!}\)